# On the Inclusion of Charge and Spin States in Cartesian Tensor Neural Network Potentials

### v1

@@ 3.3 QMspin
The QMspin dataset, introduced in Ref 24, includes molecules optimized in both singlet and triplet states. The dataset, drawn from the QM9 database, provides a comprehensive collection of around 13,000 carbene structures, with energies calculated for both singlet and triplet states. Therefore, QMspin contains approximately 26,000 data points. The singlet and triplet energy information enables the assessment of the model’s predictive performance on spin state differences, an aspect previously unaddressed due to the degeneracy of inputs, allowing us to test and validate TensorNet’s extension beyond charged states. We incorporate singlet or triplet states as $S=0$ and $S=1$, respectively, and $\lambda_{S}=\tilde{\lambda}_{S}=0.1$ for all layers.
We removed 228 data points, which corresponded to geometry files where the number of atoms in the header did not match the number of atoms with coordinate entries. Following Ref 13, we used 20,000 and 1,000 data points for training and validation, respectively, being the remaining ones used for testing. Results, found in Table 3, show that the inclusion of the spin state $S$, and therefore degeneracy breaking, improves 10-fold the accuracy of the model, achieving 43 meV error in energies, which is approximately equivalent to 1 kcal/mol, considered chemical accuracy. Furthermore, this represents an improvement of $\sim$40with respect to SpookyNet13, the only model that to the best of our knowledge has been benchmarked with the QMspin dataset.
Table 3: QMspin results. Comparison of mean absolute errors on the QMspin test set, in meV. $S$ refers to spin state, with the factor corresponding to $\lambda_{S}=\tilde{\lambda}_{S}=0.1$.
SpookyNet TensorNet TensorNet+$0.1S$
68 meV 432 meV 43 meV
@@ 4 Conclusion
In this letter, we have introduced a simple extension to TensorNet allowing it to handle charge and spin states. These attributes have typically been disregarded in state-of-the-art neural network potentials, which operate solely with atomic numbers and positions, leading to degeneracies in input space. Specifically, through a series of experiments, we have demonstrated the model’s improved accuracy and its broader applicability to a range of chemical systems.
This work opens the path to potential improvements. These results assume that additional attributes are scaled by a constant number. It would be worth exploring differentstrategies, such as making them learnable single or channel-wise scalars, potentially depending on other atomic properties such as atomic numbers. Another important aspect is that, when using partial charges, conformation-agnostic schemes have a dependence on molecular topology, which needs to be guessed from atomic coordinates, as opposed to total charge.
However, in light of the results we have so far obtained, we conclude that the approach presented in this letter represents a significant zero-cost enhancement for TensorNet, as well as an important message to take into account for future architectures.
@@ Acknowledgments
G. S. is financially supported by Generalitat de Catalunya’s Agency for Management of University and Research Grants (AGAUR) PhD grant FI-2-00587. Research reported in this publication was supported by the project PID2020-116564GB-I00, has been funded by MICIU/AEI /10.13039/501100011033, and by the National Institute of General Medical Sciences (NIGMS) of the National Institutes of Health under award number R01GM140090. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.
Appendix A Hyperparameters
All experiments were performed on two NVIDIA RTX 4090 using PyTorch Lightning’s DDP multi-GPU training protocol.
Table 4: Toy datasets hyperparameters.
@@ Appendix A Hyperparameters
Table 5: SPICE PubChem hyperparameters.
@@
Table 6: QMspin hyperparameters.

### v2

@@ 3.5 QMspin
The QMspin dataset, introduced in Ref 36, includes molecules optimized in both singlet and triplet states. The dataset, drawn from the QM9 database, provides a comprehensive collection of around 13,000 carbene structures, with energies calculated for both singlet and triplet states. Therefore, QMspin contains approximately 26,000 data points. The singlet and triplet energy information enables the assessment of the model’s predictive performance on spin state differences, an aspect previously unaddressed due to the degeneracy of inputs, allowing us to test and validate TensorNet’s extension beyond charged states. We incorporate singlet or triplet states as $S=0$ and $S=1$, respectively, and $\lambda_{S}=\tilde{\lambda}_{S}=0.1$ non-learnable for all layers.
We removed 228 data points, which corresponded to geometry files where the number of atoms in the header did not match the number of atoms with coordinate entries. Following Ref 16, we used 20k and 1k data points for training and validation, respectively, the remaining data points being used for testing. The results, found in Table 5, show that the inclusion of the spin state $S$, and therefore degeneracy breaking, improves 10-fold the accuracy of the model, achieving 43 meV error in energies, considered chemical accuracy. Furthermore, this represents an improvement of $\sim$40with respect to SpookyNet16, the only model that to the best of our knowledge has been benchmarked against the QMspin dataset. Again, SpookyNet models explicitly electrostatic, dispersion, and nuclear repulsion interactions.
Table 5: QMspin results. Comparison of mean absolute errors on the QMspin test set energies (E), in meV. $S$ refers to spin state, with the factor corresponding to $\lambda_{S}=\tilde{\lambda}_{S}=0.1$.
SpookyNet [16] TensorNet TensorNet+$0.1S$
E 68 meV 432 (19) meV 43 (2) meV
@@ 4 Conclusion
In this work, we have demonstrated and addressed a significant gap in some state-of-the-art neural network potentials, which typically disregard charge and spin states in the model’s representations. We modify TensorNet to accommodate these features with a zero-cost architectural change which effectively resolves degeneracy issues arising from the omission of these critical quantum attributes. Moreover, through experiments on well-established and custom-made benchmark datasets, we have shown that the predictive accuracy of the model is improved and that it performs on par with or outperforms existing methods that require more complex treatments of charges and spins.
Our results highlight that even a straightforward incorporation of these properties can lead to substantial improvements without the need for the introduction of additional physical terms or predicted quantities, despite facilitating and not precluding their use. Future work could explore more sophisticated strategies for scaling additional attributes, such as making them learnable scalars dependent on atomic properties.
Although this work focuses mainly on global attributes like total charge and spin, our framework mathematically supports atomic-level attributes like partial charges that could better represent local chemical environments in more complex and heterogeneous systems. However, this capability would require thorough validation and benchmarking beyond the scope of this work. Similarly, though our model effectively handles charge effects through implicit learning, it does not guarantee the correct asymptotic behavior of electrostatic interactions that physical laws demand, trading exact long-range physics for architectural simplicity while maintaining strong performance within typical molecular scales.
Overall, given the strong performance exhibited in our results, this work emphasizes the importance of incorporating charge and spin states to achieve high predictive accuracy and generality in modeling chemical systems.
@@ Acknowledgments
G. S. is financially supported by Generalitat de Catalunya’s Agency for Management of University and Research Grants (AGAUR) PhD grant FI-2-00587. A. M. is financially supported by AGAUR PhD grant 2024 FI-1-00278. Research reported in this publication was supported by the project PID2020-116564GB-I00, has been funded by MICIU/AEI /10.13039/501100011033, and by the National Institute of General Medical Sciences (NIGMS) of the National Institutes of Health under award number R01GM140090. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.
ASSOCIATED CONTENT
Supporting Information with further details on training and hyperparameters used for the different experiments presented throughout the manuscript is available.
Supplementary Information: Broadening the Scope of Neural Network Potentials through Direct Inclusion of Additional Molecular Attributes
1 Hyperparameters
We include hyperparameters used to generate the different trainings presented throughout the manuscript. Experiments were performed with the TorchMD-Net framework on two NVIDIA RTX 4090, except for Solvated Protein Fragments which was performed on four, using PyTorch Lightning’s DDP multi-GPU training protocol.
Table 6: Toy datasets training details and hyperparameters.
@@ 1 Hyperparameters
Table 7: SPICE PubChem training details and hyperparameters.
@@
Table 8: Carbon chain, silver clusters, and sodium chloride clusters training details and hyperparameters. In this case, a linear fit was performed beforehand on the training set to obtain element-wise reference energies, which were subtracted from total energy labels. For $\mathrm{Ag_{3}^{+/-}}$, y_weight was 10.0.
Parameter Value
activation silu
batch_size 4
cutoff_lower 0.0
cutoff_upper 6.0
derivative True
early_stopping_patience 300
embedding_dimension 128
equivariance_invariance_group O(3)
gradient_clipping 100
lr 1e-3
lr_factor 0.5
lr_min 1e-7
lr_patience 15
lr_warmup_steps 100
neg_dy_weight 1.0
num_layers 2
num_rbf 32
seed 1
train_size 0.9
val_size 0.1
y_weight 1.0
Table 9: Solvated protein fragments training details and hyperparameters.
Parameter Value
activation silu
batch_size 32
cutoff_lower 0.0
cutoff_upper 5.0
derivative True
early_stopping_patience 50
embedding_dimension 128
equivariance_invariance_group O(3)
gradient_clipping 100
lr 1e-3
lr_factor 0.5
lr_min 1e-7
lr_patience 10
lr_warmup_steps 0
neg_dy_weight 10.0
num_layers 2
num_rbf 32
seed 1
train_size 2560000
val_size 100000
y_weight 1.0
Table 10: QMspin training details and hyperparameters. In this case, a linear fit was performed beforehand on the training set to obtain element-wise reference energies, which were subtracted from total energy labels.